OpenStudy (lovelyharmonics):

Find the real numbers x and y that make the equation true. -3 + yi = x + 6i

OpenStudy (anonymous):

Two complex numbers are equal if and only if their real parts and equal, and their imaginary parts are equal. so for -3 + yi = x + 6i, what is the real part of the first one?

OpenStudy (lovelyharmonics):

-3 right? @Crixpack

OpenStudy (anonymous):

Correct, so what should the real part of the second complex number be?

OpenStudy (lovelyharmonics):

1x? because 6i is imaginary

OpenStudy (anonymous):

No, we have -3 + yi = x + 6i ie, we have 2 complex numbers,-3 + yi and x + 6i we want to make them equal by choosing values for x and y, Now the real part of the first one is -3 like you said, and the real part of the second one is x, which we have to choose a value now like I said, 2 complex numbers are equal if their real parts are equal and their imaginary parts are equal, so if the real part of the first one is -3, what value do we have to give to x so that it is equal to -3

OpenStudy (lovelyharmonics):

-3

OpenStudy (anonymous):

Good, so that settles x Now imaginary part of the first one is yi, and the second one is 6i, so what value do we have to give to y to make them equal

OpenStudy (lovelyharmonics):

6

OpenStudy (anonymous):

Well done :)

OpenStudy (lovelyharmonics):

thank you (: